OpenStudy (anonymous):

Determine whether the following sequence converge or diverge. Find the limit if it converges. an= cos(2pi n)/n. Okay, so seeing that this sequence is something big over small, I initially guessed that it diverges. But the answer is the limit goes to 0 and converges. I really don't see that. Shouldn't the sequence go to cos 2pi, since the n's cancel out on the top and bottom?

OpenStudy (amistre64):

\[\frac{cos(2pi*n)}{n}?\]

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

cos is the name of a function; the denominator doesnt work its way thru that to get to the other n

OpenStudy (amistre64):

cos(2pi) = 1 and all multiples of it look to be 1 as well; this gives us the sequence related to: 1/n which diverges

OpenStudy (amistre64):

the limit as n goes to 0?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so what happened to the n on top?

OpenStudy (amistre64):

you use a bit of common sense to it :) cos(2pi) = 1; and it doesnt matter how many times you hit that, your still at 1

OpenStudy (amistre64):

cos(2pi*1) = 1 cos(2pi*2) = 1 cos(2pi*3) = 1 cos(2pi*4) = 1 cos(2pi*5) = 1 cos(2pi*n) = 1

OpenStudy (anonymous):

oh, I see! thanks-you! :)

OpenStudy (amistre64):

so this acts like: 1/n, which is the harmonic series

OpenStudy (anonymous):

sequence or series? sequence converges to zero for sure. series (the sum) is divergent.

OpenStudy (amistre64):

.... yeah, that would be good to keep track of, seq or ser ;)

OpenStudy (anonymous):

so generally, something big/small diverges and small/big converges. can that always be reliable?

OpenStudy (amistre64):

depends on what your defining as converge/diverge

OpenStudy (amistre64):

\[a_n=\frac{1}{n}\] \[\lim_{n->inf} a_n=0\] the sequence itself converges

OpenStudy (amistre64):

\[a_n=\frac{1}{n}\] \[\sum_{n=1}^{inf} a_n=inf\] the series diverges